OpenStudy (anonymous):

If the perimeter of a square is one hundred forty-four miles, what is its area? twelve square miles forty-eight square miles one hundred forty-four square miles one thousand two hundred ninety-six square miles

OpenStudy (anonymous):

Perimeter of a square is = 4*a; where a is the edge of the square Given Perimeter =144 miles= 4*a therefore u can now get value of a=(144/4)=26 miles area of square= a*a Multiply and get ur answer @marrero

OpenStudy (anonymous):

dafuk did understand any thing u said

OpenStudy (anonymous):

ok where are u stuck ?

OpenStudy (anonymous):

everything

OpenStudy (anonymous):

Did u understand the formula for perimeter of square?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

U know what perimeter signifies?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

Perimeter means the sum of boundary of a figure

OpenStudy (anonymous):

As this square, it has all sides equal, so total length of boundary i.e. perimeter= a+a+a+a as it has 4 equal sides =4*a Area of square is the region covered by it= side *side=a*a=a^2

OpenStudy (anonymous):

just give me the answer

OpenStudy (anonymous):

Okay. Area= 26^2= 676 sq.mi

OpenStudy (anonymous):

Wat?

OpenStudy (anonymous):

hold on...let me write this out

OpenStudy (anonymous):

Sorry I calculate it wrong side length will be 144/4=38 and area will be 38*38=1444

OpenStudy (anonymous):

A=1296mi² A=116P2=1/16·1442=1296mi² the answer is 1296mi²